OpenStudy (anonymous):

Which compound sentence is equivalent to |x - 5| > 4? A.x - 5 > 4 or x - 5 < -4 B.-4 < x - 5 < 4 C.-8 < 2x + 3 < 8 D.x - 5 = 4 or x - 5 = -4 E 2x + 3 > 8 or 2x + 3 < -8

OpenStudy (anonymous):

D.x - 5 = 4 or x - 5 = -4

OpenStudy (lgbasallote):

show solutions @Rohangrr for the benefit of other people who wanna study this thing for references ;D

OpenStudy (anonymous):

|x - 5|=-(x-5) or (x-5) (coz |-(x-5)| =x-5 and |x-5|=x-5 put this back to the inequality, you will -(x-5)>4 or x-5>4 x-5<-4 or x-5>4